Circular No. 8288

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004Y Further to IAUC 8287, W. Li reports the LOSS discovery, on a KAIT

image taken on Feb. 12.5 UT, of a supernova (mag 18.5) located at α = 11^{h}43^{m}28^{s}.26, δ = +21^{o}40′30′′.3 (equinox 2000.0), which is 0′′.5 east and 3′′.3 north of the center of the apparent host galaxy. A. V. Filippenko, R. Chornock, and R. J. Foley report that inspection of CCD spectra (range 400–1000 nm), obtained on Feb. 14 UT with the Keck II 10-m telescope (+ ESI), shows that SN 2004Y is of type Ia, roughly 3–4 weeks past maximum brightness. The redshift of the host galaxy, measured from narrow emission lines from H II regions close to the supernova, is 0.0692.

SUPERNOVAE 2004F AND 2004X Filippenko, Chornock, and Foley add that inspection of CCD spectra,

obtained on Feb. 13 UT with the Keck I 10-m telescope (+ LRIS; range 310–920 nm) and on Feb. 14 with the Keck II 10-m telescope (+ ESI; range 400–1000 nm), reveals that SN 2004F (cf. IAUC 8271) is a peculiar variety of type IIn, resembling SN 1994W and SN 1994ak (see Fig. 14 in Filippenko 1997, ARAA 35, 309). Very narrow Hα emission (FWHM ∼ 1000 km/s) is visible, and the bottom of the corresponding P-Cyg absorption line indicates an expansion velocity of ∼ 800 km/s. Narrow absorption lines (sometimes with associated emission) of Ca II, O I, Fe II, and Na I are also visible, and they indicate comparable expansion velocities. W. Li reports that the object faded from maximum brightness quite rapidly, according to KAIT photometry.

Corrigendum. On IAUC 8287, the offset for SN 2004X in δ is 2′′.3 north (not south).

NO SUPERNOVA 2003lr Li, Filippenko, Chornock, and Foley also report that there is no super-

nova 2003lr (cf. IAUC 8266) in UGC 2904. Further inspection of the KAIT images taken on 2003 Dec. 28.23 and 2004 Jan. 10.19 UT reveals that two different minor planets were imaged at nearly the same location (7′′ positional difference): (42805) 1999 JU_{1} on Dec. 28, and (42671) 1998 HQ_{68} on Jan. 10. Remarkably, a Keck-I CCD spectrum of an object close to the putative position of the supernova, obtained on 2004 Feb. 13.32, revealed yet another minor planet, (23017) 1999 VQ_{190}. A KAIT image taken on Jan. 16.2 showed nothing at the position of 2003lr (limiting mag ∼ 19.0).

2004 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green